OpenStudy (anonymous):

Convert 1 minute and 45 seconds to minutes before you do the calculation. There are 60 seconds in a minute. Ana is composing a piece of music as background sound for a video her friends are making. The music is 1 minute and 45 seconds long, and it has 210 beats in it. Find the beats per minute rate of the music? (Beats per minute is often called tempo in music.)

OpenStudy (anonymous):

1.45?

OpenStudy (anonymous):

no @Kendal1107

OpenStudy (anonymous):

i dont know what you mean then

OpenStudy (mertsj):

You need beats over minutes.

OpenStudy (mertsj):

So how many minutes is 1 minute and 45 seconds?

OpenStudy (anonymous):

144 i think

OpenStudy (anonymous):

give me a sec

OpenStudy (mertsj):

Whose question is this?

OpenStudy (anonymous):

nope @Kendal1107

OpenStudy (mertsj):

Why isn't the asker trying to do this problem?

OpenStudy (anonymous):

well i need help step by step n idk how to set it up @Mertsj

OpenStudy (anonymous):

this question confuses me so yea...best i stay out of it

OpenStudy (mertsj):

So tell me, raevenbennett, how many minutes is 30 seconds and how do you get it?

OpenStudy (anonymous):

im really not good at math period but it almost adds up to a minute

OpenStudy (mertsj):

Let's try to learn the concept. 50 cents is how many dollars?

OpenStudy (anonymous):

2 quarters

OpenStudy (mertsj):

You must give me the answer in dollars. 50 cents is how many dollars?

OpenStudy (mertsj):

Isn't it 1/2 dollar?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Because there are 100 cents in 1 dollar so 50 cents over 100 cents is 50/100 or 1/2. Do you understand?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So since there are 60 seconds in a minute then 45 seconds is 45/60 minutes.

OpenStudy (mertsj):

Reduce that.

OpenStudy (anonymous):

okay i get that

OpenStudy (mertsj):

Reduce it.

OpenStudy (anonymous):

subtract or divide

OpenStudy (mertsj):

Reduce the fraction.

OpenStudy (mertsj):

\[\frac{45}{60}=\]

OpenStudy (anonymous):

o.75

OpenStudy (mertsj):

So 1 minute and 45 seconds = 1.75 minutes.

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So now you can find beats per minute by dividing the number of beats by the number of minutes.

OpenStudy (anonymous):

okay so how do i write my problem out

OpenStudy (mertsj):

Just put down how you changed the 1 minute and 45 seconds to minutes and then show the division problem.

OpenStudy (anonymous):

like 1/45s?

OpenStudy (mertsj):

How many minutes is 1 minute and 45 seconds?

OpenStudy (anonymous):

60?

OpenStudy (mertsj):

So 1 minute and 45 seconds = 1.75 minutes.

OpenStudy (mertsj):

You need to reread what we have covered until you understand it.

OpenStudy (anonymous):

so is it 1/45=1.75/210

OpenStudy (mertsj):

Write this down: 45 seconds = 45/60 minutes = .75 minutes. So 1 minute and 45 seconds = 1 minute + .75 minutes= 1.75 minutes 210 beats in 1.75 minutes =210 beats/1.75 minutes = 210/1.75 beats per minute = 120 beats per minute.

OpenStudy (anonymous):

ok